Entered Answer Preview Result -3.83 -3.83 incorrect The answer above is NOT correct (1 point) College students average 7.4 hours of sleep per night with a standard deviation of 40 minutes If the amount of sleep is normally distributed, what proportion of college students sleep for more than 8.5 hours? Proportion -3.83
Answer #1
for normal distribution z score =(X-μ)/σ
here mean=       μ= 7.4
std deviation   =σ= 0.6667
probability = P(X>8.5) = P(Z>1.65)= 1-P(Z<1.65)= 1-0.9505= 0.0495